A warehouse worker is pushing a 90.0 kg crate with a horizontal force of 288 N at a speed of v = 0.850 m/s across the warehouse floor. He encounters a rough horizontal section of the floor that is 0.75 m long and where the coefficient of kinetic friction between the crate and floor is 0.357.

(a)

Determine the magnitude and direction of the net force acting on the crate while it is pushed over the rough section of the floor.

magnitude Ndirection ---Select--- up down in the same direction as the motion of the crate in the opposite direction as the motion of the crate

(b)

Determine the net work done on the crate while it is pushed over the rough section of the floor.

J

(c)

Find the speed of the crate when it reaches the end of the rough surface.

m/s
Answer #1

Given:

m = 90 kg

F = 288 N

v = initial velocity = 0.85 m/s

L = length of rough surface = 0.75 m

\mu = 0.357

a) The direction of force F is in the direction of motion.

The direction of friction force exerted on the crate is opposite to the direction of motion.

As the body slides on horizontal surface, the friction force on the crate is f = \mu N = \mu mg = 0.357*90*9.81=315.2 N .

As the value of friction force is greater than the applied force, the net force is in the direction of friction force or we can say in the opposite to the direction of motion of the crate.

The magnitude of net force is

Fnet = f - F = 315.2 - 288 =27.2N [answer]

b) The net work work done on the crate is equal to W=F_{net}*L*\cos180^{\circ} [180 degree is the angle between the net force and the displacement on rough surface]

= 27.2 * 0.75 * -1

= - 20.4 J [answer]

c) Let the speed of the crate at the end of rough surface be v'.

Applying Work energy principle:

Total work done = final kinetic energy - initial kinetic energy

\Rightarrow -20.4 = \frac{1}{2}mv'^{2}-\frac{1}{2}mv^{2}

\Rightarrow \frac{1}{2}mv'^{2}=\frac{1}{2}mv^{2}-20.4

\Rightarrow v'=\sqrt{v^{2}-\frac{20.4*2}{m}}

\Rightarrow v'=\sqrt{0.85^{2}-\frac{20.4*2}{90}}=0.52 m/s [answer]